OpenStudy (anonymous):

what is the inverse of 7x-5

OpenStudy (stefrheart):

What is the definition of an inverse?

OpenStudy (anonymous):

\[\frac{ 1 }{ 7x-5 }\]

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

x=7y-5 7y=x+5 f^-1(x)=(x+5)/7